Suppose that 97% of the population is not lactose intolerant. If you were to randomly select...

  1. Suppose that 97% of the population is not lactose intolerant.
  1. If you were to randomly select 3 people at random, what is the probability that all of them are not lactose intolerant? Assume that each selection is made with replacement.
  2. If you were to randomly select 3 people at random, what is the probability that at least one of them is lactose intolerant? Assume that each selection is made with replacement.
  3. If you were to randomly select 3 people at random, what is the probability that only the third person selected will be lactose intolerant? Assume that each selection is made with replacement.
Answer #1

a)

0.97^3 = 0.91267

b)

P(atleast 1 is latose intollerant) = 1- P(all 3 not latose intolerant) =1 - 0.91267 = 0.08733

c)

P(x=1) = C(3,1)*(0.97^2)*(0.03) = 3*(0.97^2)*(0.03) = 0.08468